### v1

# AugGen: Augmenting Task-Based Learning in Professional Creative Software with LLM-Generated Scaffolded UIs

Abstract.
Professional creative software often presents steep learning curves due to complex interfaces, lack of structured task-aware guidance, and unfamiliar domain terminology. To address these challenges and augment user learning experience, we introduce AugGen, a method for generating scaffolded user interfaces that simplify interface complexity and support task-based learning. With the user’s task, our method surfaces task-relevant tools to reduce distracting features, organizes the tools around task workflow stages to offer execution guidance, connects tools with domain concepts to foster learning engagement, and progressively discloses advanced features to manage learning progress. To evaluate the method, we used our LLM-assisted pipeline to generate two task-specific scaffolded UIs and deployed them in Blender, our professional 3D modeling testbed. We invited both beginner (N=32) and expert (N=8) users to evaluate our implemented interfaces. Results show that the scaffolded interfaces significantly reduced user-perceived task load, enhanced task performance via embedded guidance, and augmented concept learning during task execution.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06
Figure 1. With AugGen, users can generate scaffolded user interfaces (UIs) to augment software and domain concept learning while executing their tasks in professional creative software. Taking a task description as input, AugGen automatically analyzes user task to decompose the task into workflow stages and extract relevant tools and domain concepts. The functionality of selected tools are implemented through programming and presented to the user as UI controls on the scaffolded interface. The code to implement functionality and UI controls are generated by LLMs and executed in the software for user interaction. Users can interact with the scaffolded interface with task-relevant tools structured around the workflow stages and learn domain concepts through labeled tooltips during task execution.
1. Introduction
Professional creative software, such as Photoshop (photoshop), Illustrator (illustrator), and Blender (blender), offers significant skill augmentation and forms the backbone of many creative and design industries. However, this skill comes at a cost of increased learning complexity. Prior work has identified that a major complexity arises from the user interface (UI), where newcomers often encounter hundreds of icons, menus, and specialized tools (chilana2018supporting). Learners frequently struggle to find the appropriate functions or tools and understand domain-specific terminology (novick2009micro; kiani2019beyond; lee2010usability). Compounding these issues is a lack of built-in structured guidance, which pushes users toward external resources, such as video tutorials, forums, and increasingly, large language models (LLMs), to navigate the UI and meet their task and learning goals. The continual toggling between disparate help sources disrupts workflow and frequently fails to deliver contextually appropriate support, thereby hindering self-directed learning and exploration (rieman1996field; chilana2012lemonaid; chilana2018supporting; khurana2024and).
Multiple approaches have attempted to mitigate these challenges. One common strategy has been to provide simplified software alternatives, such as dedicated beginner modes (mcgrenere2002evaluation), separate novice-oriented applications with limited functionality (tinkercad; illustratoripad), or AI-generated task-specific interfaces (vaithilingam2024dynavis). While these alternatives reduce initial complexity, they often embody a simplicity-power tradeoff (uiandvis1999hearst), where users may plateau in simplified environments (cockburn2014supporting), or face a context switch when transitioning to the professional suite, which negates prior learning (nicol2024psychology; li2022design).
Another line of work has focused on improving in-context assistance. Efforts range from static tooltips (grossman2010toolclips) and context-sensitive help (kelleher2005stencils) to more sophisticated systems that embed external resources directly into the application context (zhong2021helpviz). However, these approaches struggle with accurately inferring user goals, potentially leading to intrusive or irrelevant help (brdnik2022intelligent), and users may still find it challenging to map the provided information onto the complex software interface (kosch2023survey).
More recently, the advancement of LLMs has led to the emergence of in-application software copilots that can process natural language queries, monitor visual context, or automate tasks (firefly; krol2025copilotvision; khurana2025me). While powerful for automating well-defined operations, current copilots struggle with complex, creative, or exploratory tasks (khurana2024and; qiao2024benchmarking). Furthermore, prior work has found that excessive automation or guidance by these copilots can reduce engagement (kosmyna2025your), diminish agency (li2024user; sellen2024rise), reduce the user’s ability to learn underlying concepts (chen2023understanding; vorvoreanu2025fostering), and contradict the well-established maxim of learning by doing (carroll1984training; rieman1996field). Lastly, excessive automation and prescriptive guidance from copilots can lead users to adopt a passive, instruction-following mode, akin to the studied “GPS effect” (ruginski2019gps), instead of learning.
Despite substantial prior work, existing systems still fall short of holistically reducing UI complexity, offering in-context structured guidance, and helping users interpret domain-specific terminology in ways that integrate learning into task execution. Motivated by instructional-scaffolding research where temporary support helps learners build competence and gradually work independently (belland2017instructional), we introduce AugGen, a method that generates scaffolded interfaces using the underlying native software’s UI elements and operations. As shown in Figure 1, we demonstrate our LLM-based approach to generatively surface task-relevant tools from the underlying software and present them in a simplified layout tailored to the user’s current task. The resulting scaffolded UI progressively expands in complexity as the user gains proficiency, links interface elements with domain concepts to support concept learning, and guides the user through task workflows while remaining grounded in the native software environment.
We apply our method to build two scaffolded interfaces in Blender (blender), a professional open-source 3D creation environment used as our testbed. We evaluate the two interfaces, one for 3D modeling and one for animation, with beginner (N=32) and expert (N=8) users. Results show that compared to the default Blender interface, the scaffolded UI has significantly improved task performance for both groups, as well as augmented concept learning for beginners. Based on a discussion of these findings, we present insights for future research on using our method to create scaffolded interfaces for instructional use cases to support both educators and learners, and to allow personalized workflows and cross-platform synchronization for end users. In summary, we make the following contributions:
  • We propose a scaffolded interface generation method to reduce UI complexity of professional creative software, streamline task guidance through UI configuration, and augment domain concept learning during task execution.
  • We present a technical pipeline that applies the method to generate scaffolded interfaces in professional creative software based on LLM-assisted UI generation.
  • We conduct a user study with both novices and experts to demonstrate the effectiveness of the scaffolded interface generation method for task execution and concept learning support.
2. Related Work
Our work builds upon prior research to address learning related challenges in professional creative software. Figure 2 presents a visual summary.
Figure 2. Challenges in professional creative software (C1–C3), prior research to tackle these challenges, and AugGen’s focus.
Challenges in professional creative software. Professional creative software has long been recognized as having a steep learning curve. A primary challenge is the complexity of the interface which presents numerous icons, tool palettes, and menus (chilana2018supporting). Prior work has identified that beginners frequently have trouble finding tools and determining the operational sequence for their tasks (novick2009micro; lee2010usability; kiani2019beyond). Another challenge is that the software provides limited guidance on how to proceed with user tasks. To address these challenges, users often turn to learning resources which are mainly external, e.g., formal manuals (novick2006don), tutorial videos (lafreniere2013community), blogs, and dedicated forums, which disrupt task workflows and create a high barrier to self-taught exploration (rieman1996field; chilana2012lemonaid; chilana2018supporting; khurana2024and). Additionally, professional creative software is designed for domain experts and its interfaces are often saturated with complex domain terminology, which raises difficulty and steepens the learning curve for beginners (popovic2000expert). In summary, the challenges of professional creative software that obstruct the learning experience include C1: complex UIs with extensive features; C2: limited context-aware, built-in guidance; C3: unfamiliar domain concepts.
Figure 3. AI copilots in Blender that offer step-by-step task instructions (left) (blender_copilot_oai) or automatically execute a task (right) (blender_copilot). The copilot’s instructions overlook interface cues and operational context. By assuming familiarity with Blender’s modes, shortcuts, and domain concepts, the guidance offers an incomplete procedural scaffold for beginner users.
Simplified alternatives to professional creative software. To tackle these challenges, one approach has been to build software with simplified functionality for beginners, which addresses C1. For example, Adobe Express (adobeexpress), Canva (canva), OpenShot (openshot), TinkerCAD (tinkercad), and Illustrator on iPad (illustratoripad) target newcomers with simpler UIs and basic operations. In addition, recent work has explored natural language interfaces (generativefill; qin2024instructvid2vid; poole2022dreamfusion; liu2024dancegen) or LLM-generated interfaces (vaithilingam2024dynavis; liu2025crowdgenui), with simple interaction modalities like text or on-demand UI controls. These alternatives embody the classic simplicity-power tradeoff (uiandvis1999hearst), where simplified interfaces are less intimidating for novices, but they often lack the breadth of functionality. To balance this tradeoff, prior research has explored the transition between multiple interface modes, such as multi-layer interfaces (shneiderman2002promoting; mcgrenere2002evaluation) and workflow documentation interfaces (grossman2010chronicle).
However, the gap between simple and advanced interfaces may lead to users getting stuck in the beginner mode if the two modes are too different (cockburn2014supporting), or getting overwhelmed if advancing to complex modes is poorly managed (nicol2024psychology). Another issue is that context switching from novice to professional software requires users to relearn workflows because the interfaces differ. Such discontinuity can void some of the initial learning done in the simplified application (li2022design). Our work aims to address these limitations by embedding a scaffolded interface into professional creative software to reinforce connection with the native software and reduce discontinuity from platform switching. To manage UI complexity and help users move from basic to advanced, we offer user-selectable complexity levels that let them gradually take on more advanced tools as they learn the underlying concepts and work through their tasks.
In-context assistance for professional creative software. To address C2, research efforts have explored in-context assistance in professional creative software. Unlike external resources, such assistance aims to provide support at the user’s point of need, thus reducing disruption. For instance, previous work has designed elements, such as tooltips or help buttons, that are linked to relevant sections of a manual (farkas1993role; kelleher2005stencils; grossman2010toolclips; fourney2014intertwine; photoshop). While integrated into the application, these elements are static and often lack user context sensitivity. Recognizing the limitations of static help, researchers have developed context-aware approaches. One thread of work explores intelligent user assistance, such as using Bayesian probability networks to model user goals and offer help without explicit user requests (li2005active; horvitz2013lumiere). Another line of work focuses on improving contextual assistance delivery by bringing external resources into the application context, such as automatically presenting relevant web pages, documentation sections, forums, or tutorial videos based on the user’s current activity (chilana2012lemonaid; fraser2019replay; matejka2011ambient; matejka2011ip; ekstrand2011searching; zhong2021helpviz; yang2022softvideo; yang2024aqua).
Nonetheless, existing in-context assistance faces challenges in accurately inferring user goals and risks providing irrelevant or poorly timed assistance (brdnik2022intelligent). In addition, while contextual assistance delivery reduces the need for explicit searching, users can still struggle to identify the most relevant information (lim2012improving) and map the provided guidance (textual steps or visual cues) onto the complex interfaces (kosch2023survey). We tackle these challenges by building a task-aware interface that surfaces the relevant tools from the underlying software to avoid contextually irrelevant support. Rather than asking users to hypothesize instructions and navigate a complex UI, our scaffolded interface embeds the guidance by grouping tools into workflow stages, aiming for a clear task-tool alignment.
In-application software copilots. The recent advancement of LLMs has reshaped the landscape of in-context assistance, leading to the emergence of in-application AI copilots. These copilots can process natural language queries, offer instructions, and automate software tasks, e.g., copilots in Blender as shown in Figure 3. These copilots can also use multimodal information, such as screen monitoring, to understand user context and offer guidance, e.g., Microsoft Copilot Vision (krol2025copilotvision; cunningham2025copilotvision) and Google AI Studio (google2025aistudio). More recently, the copilots have unlocked the potential for on-demand UI generation (e.g., Gemini 3 (gemimi3)), which could be used to generate dynamic UIs to assist tasks based on user needs.
Although these copilots are helpful to guide or automate tasks, excessive guidance and task automation can diminish the user’s sense of ownership over the task (endsley2017here; heer2019agency; li2024user; sellen2024rise) and reduce cognitive engagement in learning (kosmyna2025your). These copilots surface foundational concepts only incidentally, hindering a user’s conceptual understanding (C3) and running counter to the principle of learning by doing (carroll1984training; rieman1996field; novick2009micro; kiani2019beyond). The copilots might also suggest tools or UI elements that do not exist in the software or are not directly available, making the learning experience even more challenging. Although on-demand UIs can be generated by copilots to assist task execution, prior research has found that constantly changing UIs can increase cognitive load for the user, especially during longer task sessions (vaithilingam2024dynavis). To address these challenges and support durable learning, we embed explicit instructional scaffolds into the UI, aiming to engage the user in tasks. We bind each task-related action to its underlying concept to support learning while executing tasks. Furthermore, to reduce the cognitive load caused by dynamic UIs that may hinder learning, we offer users ready-to-use scaffolded UIs tailored to their tasks.
3. AugGen
In this section, we present our method and implementation to generate task-specific scaffolded interfaces for professional creative software by surfacing relevant tools, embedding domain concepts, and mapping the scaffolded interface back to the native software. The method is grounded in principles of task-centric design, adaptive interfaces, and scaffolded learning, and implemented as a multi-stage pipeline based on LLM-assisted UI generation.
3.1. Method
Task-specific tool mapping. The method begins by extracting a set of tools required for the user’s task. Hidden or nested functions are surfaced to reduce learning barriers and interface complexity (C1). This is motivated by task-centric UI design principles (lafreniere2014task) to manage interface complexity and reduce user cognitive load (sweller2011cognitive).
Scaffolded interface construction. The method constructs an interface that organizes selected tools into workflow stages to offer task guidance and eliminate the need for users to switch between external tutorials and the software interface (C2). This is grounded in adaptive UI design principles (knowledge_navigator; jameson2007adaptive) to offer in-context, task-aware guidance through UI configuration.
Progressive disclosure. The scaffolded interface supports user-selectable levels of complexity, which enables beginners to start with fundamental functions and expand the interface as they gain proficiency. This aligns with the training-wheels interface model (carroll1984training) and intermodal guidelines (cockburn2014supporting) to manage the learning process and encourage exploration with user control.
Concept labeling. Each tool is labeled with domain terminology to connect actions, concepts, and workflow steps, supporting concept learning during task execution (C3). This is motivated by prior research on learning by doing (rieman1996field; novick2009micro) and vocabulary extension (cockburn2014supporting) to explicitly reinforce learning engagement while performing tasks.
Native-UI mapping. To prevent over-reliance on the scaffolded UI, each tool is linked to its location in the native environment through shortcut labels, menu paths, and interaction cues, which facilitates gradual transition toward native software fluency.
3.2. Implementation
3.2.1. Technical Pipeline
Here we present a technical pipeline to implement the scaffolded interface. In this work, we demonstrate an example that implements the scaffolded interface in Blender (version 3.6), a professional 3D creative software tool. The implementation pipeline involves using an LLM (GPT-4o) to assist with user task analysis, task-relevant tool selection, and UI code generation. This section outlines the technical pipeline and the complete implementation details for reproducibility can be found in Appendix A. Figure 4 shows this pipeline.
Figure 4. Technical implementation pipeline built upon the AugGen method. (1) The pipeline takes a user’s task description in natural language (e.g., perform UV unwrapping) that informs LLM-assisted task analysis to reason workflow stages and relevant domain concepts (e.g., marking seams, unwrapping, checking and visualization). (2) These workflow stages subsequently guide LLM-supported tool selection and complexity assessment, resulting in a candidate tool list (e.g., tools to mark seams and unwrap UV map). (3) The tool list is fed into LLM-based code generation for the UI implementation. (4) The code is executed to generate a scaffolded interface in professional creative software (e.g., custom panel in Blender). Users (e.g., UI designers, software educators, end users) are allowed to step in to refine LLM-generated artifacts (e.g., workflow, tool list, code) to adjust the scaffolded interface.:
Workflow analysis and concept identification. The technical pipeline starts with taking a user’s task description in natural language as the input. AugGen then assembles the workflow analysis and concept identification prompt (Appendix A.1) that guides the LLM to decompose the user task into sub-tasks, forming the task workflow, and identify relevant domain concepts, which are associated with each sub-task. In this process, the LLM needs to analyze the required steps to perform the user task and classify these steps into workflow stages, as well as extract relevant domain concepts and explain them in a straightforward, concise description.
Tool selection and complexity assessment. With the workflow stages and relevant concepts, AugGen selects needed tools for each stage and maps them to the software functionality. For each workflow stage, AugGen uses the tool selection prompt with reference to the software manuals for the LLM to select tools, e.g., keyboard and mouse operations and UI interactions. As part of this process, AugGen assembles the complexity assessment prompt to instruct the LLM to suggest a complexity level for each tool (e.g., basic, intermediate, advanced) based on its functionality from software manuals (e.g., manuals in (blendermanuals)), and add these complexity levels to the tool description. Prompts for tool selection and complexity assessment are available in Appendix A.2.
UI code generation and tool labeling. With the selected tools, AugGen generates the code to implement each tool’s functionality and renders the tools in a scaffolded interface. Specifically, AugGen first assembles the functionality code generation prompt for the LLM to generate the code that executes each tool’s functionality by referring to the software APIs (e.g., software APIs in (blenderapidocs)). Essentially, each keyboard and mouse operation or UI interaction is translated into an executable functionally-equivalent code block. Next, AugGen uses the UI code generation prompt to guide the LLM to generate the code that converts the functionality code execution into UI interaction, e.g., button click or drop-down selection, such that users can initiate the execution of each tool through the scaffolded interface. In the UI code generation process, AugGen provides the LLM with an example code that outlines the interface layout to structure tools based on their corresponding workflow stages and groups tools according to their complexity levels. Lastly, AugGen assembles the tool labeling prompt to ask the LLM to add concept description as tooltips for each tool. Prompts for code generation are available in Appendix A.3.
Code execution and UI testing. In the last stage, AugGen executes the generated UI code in the professional creative software to display the scaffolded interface (e.g., a custom panel in Blender). With this interface, UI designers, software educators, and end users, can test and evaluate the UI or make optional adjustments. They can adjust the intermediate artifacts offered to them in text, including the workflow stages, domain concepts, and tool selections, as well as the generated UI code for UI customization. These adjustments lead to iterative executions of the technical pipeline for the user to obtain updated scaffolded interfaces. Each execution takes 1-2 minutes.
3.2.2. Generalizability
This technical pipeline can be used to accommodate different user tasks in the same and different software tools. Broader tasks are achievable by substituting a new task description, in natural language, as input to this pipeline. The pipeline will then analyze the task workflow, extract relevant domain concepts, select tools, and generate the UI code to create new scaffolded interfaces. Beyond a single software tool, this pipeline can be used to generate scaffolded interfaces in other software that provides a manual of its functionality for workflow analysis and tool selection, and programming APIs for UI code generation. Example software include Photoshop (photoshopapi), Maya (maya_api), AutoCAD (autocad_api), Unity (unity_api), and Grasshopper (grasshopperapi).
4. Illustrative Scenarios
To illustrate the user experience in using our scaffolded interfaces, we walk readers through how a user can complete a 3D modeling task and learn relevant domain concepts. We compare the experience with performing the same task using the default interface of a professional creative software tool (Blender).
Task introduction. UV unwrapping is a process for applying 2D textures to a 3D model’s surface and requires domain concept understanding. It involves cutting and unfolding the 3D model into a flat 2D representation called a UV map. Each point on this 2D map, defined by U and V coordinates, corresponds to a vertex on the 3D model, allowing textures to be projected onto its surface. The major process includes defining seams—edges on the 3D model where the mesh will be split, and then executing an unwrap operation to generate the 2D layout of UV islands. Thereafter, these islands are arranged and optimized in the 2D space, and finally, the layout is checked to reduce distortions or overlaps.
Experience using the default UI. Performing this task in the default UI (Figure 5) requires the user to navigate different menus and modes. The user enters “Edit Mode”, switches to “Edge Selection” mode, and then starts selecting edges on the 3D mesh. After selecting the edges, the user marks them as seams in a dedicated UV menu via “Ctrl+E”. Following it, they “right-click” within the editing region to open an unwrapping operator to unwrap the 3D mesh along the marked seams. The resulting UVs are viewed and refined in a separate UV Editor using tools from a toolbar.
During the process, the user spends significant amount of time following tutorials and familiarizing themselves with the operations and UI, such that they can locate and use tools through keyboard shortcuts, right clicks, and the UV Editor’s toolbar to complete the task. In addition, domain concepts relevant to their task are not integrated into their task execution but instead requiring additional learning effort by searching online or taking courses.
Figure 5. UV unwrapping using the default UI of Blender. Tools are located sparsely and sometimes hidden in keyboard shortcuts or mouse right-clicks. Domain concepts like “seams”, “unwrap”, and “UV islands” are opaque to users with limited explanation.:
Experience using the scaffolded UI. Using our scaffolded interface, as shown in Figure 8, to work on this task, the user starts at the Basic complexity level (Figure 6) and sees a structured task workflow. They first work through the Marking Seams stage. All the tools needed to select edges and mark seams are organized under this stage. Following the suggested flow, they move to Unwrapping & Editing and use the tools to unwrap the UV map and perform a simple visual check using tools in Checking & Visualization. Seeking to improve this initial result, they switch the interface’s complexity level to Intermediate (Figure 7) to utilize newly available seam selection aids and refine the UV layout. For final polishing and access to the complete toolset, they select the Advanced level (Figure 8) to reveal specialized tools. At this level, they can employ advanced operators for accurate adjustments and use analysis features to minimize distortion and verify the UV map quality.
Figure 6. UV unwrapping Basic level of the scaffolded UI.:
Figure 7. UV unwrapping Intermediate level of the scaffolded UI.:
While working on the task, the user can reveal relevant domain concepts by hovering over each tool to see the tooltips. For instance, “Mark Seam: Where to ‘cut’ the 3D model’s surface so it can be unfolded into a flat 2D layout”, “Unwrap: Flattens the 3D model’s surface into 2D space based on the marked seams”, and “UV Islands: Makes UV islands (separate unwrapped parts) have the same relative size so textures don’t look distorted”. They can also see the keyboard shortcuts, mouse clicks, and native locations of each tool in its label to get familiar with the native software environment.
Throughout the process, the user is guided by the task workflow indicated by the interface without the need to switch back and forth between tutorials and their task. They can find task-relevant tools grouped under the corresponding workflow stage with less struggle remembering tool locations or the dedicated keyboard shortcuts and mouse clicks. When working on their task, they can control the tool complexity from basic to advanced based on their expertise and progress. They also get the chance to link their operations to relevant domain concepts and learn these concepts during task execution. Although the scaffolded UI is shown on a separate custom panel, they can connect each tool with the default software’s menu, keyboard shortcuts, and mouse operations through tool labels.
Figure 8. UV unwrapping using the scaffolded UI. Task-relevant tools are selected and grouped based on the workflow stages. Users can select complexity levels from basic, intermediate, to advanced for progressive tool disclosure. Hovering over each tool shows a tooltip explaining the related domain concepts. To connect the scaffolded interface with the native software environment, keyboard shortcuts, mouse clicks, and the location of each tool in the menus or toolbars are labeled.:
5. User Study Design
To evaluate our scaffolded interfaces, here we introduce the user study design with beginner (Study 1; Section 5.1) and expert users (Study 2; Section 5.2).
5.1. Study 1: Beginner Users
From the study with beginners, we explore the following research questions (RQs) to evaluate the effectiveness of the scaffolded interface compared to the baseline (default Blender interface) regarding task performance, concept learning, and software experience.
  • RQ1: How does the scaffolded UI influence task performance? This question explores whether task-aware interfaces with relevant tools impact user-perceived task load, task performance, task completion time (efficiency), and user confidence.
  • RQ2: How does the scaffolded UI affect the learning of domain concepts and software operations? This question evaluates how progressive tool disclosure, concept-integrated UI organization, and bridging to native software help users understand domain concepts, connect concepts to tools, and learn the native software operations.
  • RQ3: How does the user experience differ with the scaffolded UI? This question investigates the overall user experience, including user-perceived ease of use, sense of support vs. constraint, entry barriers, and reliance on external help and documentation.
Experiment design. We adopted a between-subjects design where each participant was randomly assigned to one of two interface conditions (default Blender interface or our scaffolded interface) and to one of two task conditions. Task 1 asked participants to perform UV unwrapping on a cube, a beginner-level texturing task, while Task 2 involved constructing a human walk cycle, a basic animation task. We chose these tasks because they represent two main task pipelines of 3D modeling and animation and require domain concept understanding. These tasks also accommodate adjustable complexity levels for us to evaluate with both beginner and expert users (Section 5.2). We provided all participants with the same cube for Task 1 and the same rigged human character ready for animation for Task 2 to ensure experiment consistency. The interfaces for Tasks 1 and 2 are shown in Figure 8 and Appendix B.1, respectively. To ensure consistent study instruction, we provided a study guide (Appendix B.2) that introduced the basics of UV unwrapping and animation, as well as publicly available tutorial videos, with durations of 4.5 minutes for Task 1 and 7 minutes for Task 2. Task 2 was more complex than Task 1 due to the longer completion time and more required steps.
Participant recruitment. We recruited 32 English-speaking beginners (11 female, 19 male, 2 non-binary; mean age = 27.31, SD = 7.34) with no or limited 3D modeling and animation experience via Prolific (100% approval rate) (prolific). Participants were randomly assigned in a fully balanced design to one of two interface conditions, (1) Baseline (default Blender interface), or (2) Ours (scaffolded interface), to complete either Task 1 or Task 2, resulting in eight participants per task-interface combination.
Measurements. We used four questionnaires with 7-point Likert scale responses. For RQ1, we used the NASA Task Load Index (TLX) (hart2006nasa) and a custom task performance questionnaire. For RQ2 and RQ3, we used custom questionnaires on concept learning and user interface experience. These custom questionnaires were designed based on broadly used standard questionnaires and an instructional model, including Software Usability Measurement Inventory (SUMI) (kirakowski1996software), Instructional Materials Motivation Survey (IMMS) (keller2010imms), and ARCS model of motivation (keller2009motivational). All the questionnaires are available in Appendix B.3.
Procedure. We invited eligible participants to a Zoom meeting. Before the study started, we emailed a consent form (IRB protocol # anonymized), which outlined the study goal, procedure, duration, and data collection. After providing informed verbal consent, participants filled out a pre-study demographics questionnaire ($\sim$5 minutes). We then shared a study guide via a Google Slides link in Zoom chat, which included task instructions and tutorials. Once participants confirmed they had read the guide and accurately described the task, we shared our desktop screen, gave them remote control, and asked them to complete their assigned task. With consent, we recorded the desktop screen to record their task actions. Participants were encouraged to think aloud, and we took notes when observing them. Based on a pilot study with two beginners that helped us estimate task completion time, we gave the participants 20 minutes for Task 1 and 30 minutes for Task 2. After the task, participants were invited to complete the post-study questionnaires ($\sim$10 minutes). The study concluded with a payment of 10 USD to participants who worked on Task 1 for completing the $\sim$35 minute study and 15 USD to participants who worked on Task 2 for completing the $\sim$45 minute study.
Data analysis. To evaluate the effects of interface and task on participant responses, we conducted Type II two-way analyses of variance (ANOVA) (seabold2010statsmodels) with Interface (Ours vs. Baseline) and Task (Task 1 vs. Task 2) as between-subjects factors to examine the main and interaction effects on our measurements. Following significant ANOVA effects, we performed post-hoc pairwise comparisons using Bonferroni-adjusted t-tests (Vallat2018pingouin) to identify specific differences.
5.2. Study 2: Expert Users
For the study with experts, we kept RQ1–RQ3 for a comparative analysis across beginner and expert users. In addition, two further research questions were formulated to explore the impact on skill acquisition and method generalization.
  • RQ4: How can the scaffolded UI potentially shape long-term skill development? This question invites expert perspectives on whether the scaffolded interface can help refine skills, adapt tools to new tasks, and strengthen concept understanding for solving new problems.
  • RQ5: What is the potential impact of applying the scaffolded UI generation method to other software? This question asks for insights from experts familiar with additional professional creative software beyond Blender on the benefits and challenges of using scaffolded interface in different platforms.
Experiment design. Each participant was assigned to one of two task conditions. Task 1 involved performing UV unwrapping on a 3D monkey head (Figure 8), an advanced texturing task, while Task 2 required constructing a human walk cycle that reflected personality, an open-ended animation task. We provided the 3D model for Task 1 and the rigged human character for Task 2 for experiment consistency. Task assignment was based on participant prior experience: only those with 3D model texturing experience were assigned to Task 1, and only experienced animators were assigned to Task 2. This allowed us to gather feedback grounded in a participant’s prior expertise to enable comparisons between their experience with the Blender interface and the scaffolded interface.
Participant recruitment. We recruited eight experts (3 female, 5 male; mean age = 28.87, SD = 3.66) via Prolific (prolific), requiring at least one year of Blender experience and preferring additional proficiency in tools like Maya or 3ds Max. All other criteria for gender, age, language, and Prolific history matched those of Study 1. Participants (S2-P1 to S2-P8) were each assigned to either Task 1 or Task 2 and completed their task using the scaffolded interface, with four experts per task. Detailed demographics appear in Appendix C.
Measurements. To evaluate RQs 1, 2, and 3, we used the same questionnaires as Study 1 and adapted the items to preference-based questions: each item was rated on a 7-point Likert scale (1 = strongly prefer our interface, 7 = strongly prefer the baseline interface). Additionally, we collected qualitative feedback by asking participants to elaborate on their experience with our interface regarding task performance, concept learning, and interface experience. For RQs 4 and 5, we gathered user qualitative data regarding skill development and using the scaffolded interface in other software tools. Refer to Appendix B.4 for the open-ended questions for qualitative data collection.
Procedure. We invited eligible participants to our study via Zoom. The pre-study session involving the consent and pre-study questionnaire had the same procedure as Study 1. Following this, we offered a brief introduction to the user’s task ($\sim$5 minutes). After participants verbally confirmed that they were clear about the task requirements, they used the scaffolded interface to work on the task for 15 minutes through remote desktop control. We recorded the desktop screen with participant consent and took notes on their actions. After the task, participants completed the post-study questionnaires and answered the open-ended questions ($\sim$10 minutes). The study concluded with a payment of 10 USD for completing the 30-minute study.
Data analysis. The qualitative data was analyzed using thematic analysis (braun2006using) with a deductive strategy. The authors reviewed the participant responses, coded the data to extract relevant information centered around the research questions, refined the codes, and categorized themes.
5.3. Study Design Remarks
Since our study with beginners aimed to explore their learning experience using scaffolded UIs in a complex creative tool rather than making strong causal claims, not a large sample size was the main priority. We used a between-subjects design to avoid cross-condition contamination during learning and skill acquisition. In the study with experts, we prioritized qualitative depth over large-scale inference, and the eight users frequently interacted with Blender and were able to articulate pros and cons of our scaffolded interface clearly. Also, we observed signs of saturation during thematic analysis of the qualitative data.
6. User Study Results
In this section, we summarize the user study findings with beginner (Section 6.1) and expert users (Section 6.2). The findings are organized in response to our research questions (as introduced in Sections 5.1 and 5.2).
6.1. Study 1: Beginner Users
6.1.1. Questionnaire Responses
Figure 9 shows the participant reported responses to the questionnaires. Asterisks in the figure indicate statistical significance derived from post-hoc pairwise comparisons between Ours (scaffolded) and Baseline (default Blender) interfaces for each task and measure. The complete two-way ANOVA results are reported in Appendix D. From the results, we see that our scaffolded interfaces have significantly improved task performance, concept learning, and user experience compared to the default Blender interface.
(a) Task load.
(b) Task performance.
(c) Concept learning.
(d) User interface experience.
Figure 9. Study 1 participant responses using Ours (scaffolded) and Baseline (default Blender) interfaces for Tasks 1 and 2 (lower is better). Asterisks denote significant differences in post-hoc pairwise comparisons between interfaces within each measure (*: $p<.05$, **: $p<.01$, and ***: $p<.001$).:
6.1.2. Task Completion Time
All participants successfully completed their tasks and generally spent less time using our scaffolded interface compared to the default Blender interface: $\bar{t}_{Ours(Task1)}=12.69\pm 1.49$ mins vs. $\bar{t}_{Baseline(Task1)}=15.08\pm 2.34$ mins; $\bar{t}_{Ours(Task2)}=20.39\pm 2.55$ mins vs. $\bar{t}_{Baseline(Task2)}=24.95\pm 3.46$ mins. Significant main effects of Interface ($F_{(1,30)}=14.78,p<.001,\eta^{2}_{p}=.33$) and Task ($F_{(1,30)}=94.71,p<.001,\eta^{2}_{p}=.76$) were observed. The interaction effect was not significant. Post-hoc analysis revealed no significant difference in completion time in Task 1, but users with our interface were significantly faster in Task 2 ($p<.05$), demonstrating the efficiency advantage of our scaffolded interface under higher task demands.
6.1.3. User Behavior
We observed that participants using our scaffolded interface (experiment group) progressively reduced their reliance on tutorial videos and increasingly navigated via the interface’s inherent workflow structure, echoed by participant self-reported improved workflow clarity, tool relevance, and reduced external help requirement. In contrast, participants using the default Blender interface (control group) consistently referred to the videos throughout the tasks. Consequently, the experiment group generally completed tasks faster than the control group. Additionally, when some tools were not mentioned in the tutorials but presented in our scaffolded interfaces, the experiment group explored them during or after their tasks, as echoed by participant self-reported tool and concept exploration. Conversely, the control group rarely explored beyond the tools covered in tutorials.
6.2. Study 2: Expert Users
6.2.1. Questionnaire Responses
Figure 10 shows the participant reported preference scores for Ours (scaffolded) or Baseline (default Blender) interfaces on the measures of task performance, concept learning, and user interface experience. They had a clear preference for our interface across all dimensions regarding task performance. For concept learning, participants generally perceived our interface as more supportive for understanding concepts, relating tools to concepts and workflows, and motivating learning. However, they had mixed opinions on concept exploration (4 favored our interface, 1 was neutral, 3 favored the baseline). For the overall interface experience, consistent preference for our interface was observed except for tool exploration (5 participants preferred our interface, 3 favored the baseline).
Figure 10. Study 2 participant preferences for Ours (scaffolded) or Baseline (default Blender) interfaces regarding task performance, concept learning, and user interface experience.:
6.2.2. Qualitative Results
Here we present the themes derived from qualitative data we obtained from the expert users.
Tradeoffs of supporting task performance and allowing workflow flexibility (RQ1). Participants generally appreciated task performance support by surfacing task-relevant tools (S2-P1, S2-P2, S2-P7, S2-P8). The scaffolded interface led to higher task efficiency and user focus, as S2-P2 said that, “I actually finished the UV task faster than expected.” However, participants also expressed concerns that the system’s assumptions sometimes misaligned with their preferred workflows. Some found the interface limiting when it failed to surface user-expected tools (S2-P4). Others perceived the guided structure forced a specific process, which may reduce workflow flexibility (S2-P6). These findings indicate the need for more user customization to support expert users who prefer flexible or iterative workflows.
Learning and exploration opportunities during task performance (RQ2). Participants liked our interface for its ability to clarify the task-related concepts, which can benefit educational or onboarding contexts (S2-P1, S2-P3). They also highlighted the learning benefits, saying that, “it’s not just clicking through menus but also teaching you why they matter” (S2-P2). In addition, progressive tool disclosure was appreciated as a way to guide structured concept learning. For instance, S2-P3 noted that, “it shows graph editing after you’re done with key posing, and this can help avoid overwhelming learners.” Furthermore, participants remarked that showing tools adjacent to the immediate task was helpful to motivate learning. As noted by S2-P5, “It prompted me to notice things like secondary motion I’d normally skip past,” indicating that the interface enables the discovery of related concepts and tools and encourages user exploration.
Gains of simplicity and focus vs. loss of familiarity (RQ3). Participants responded positively to the overall experience since the scaffolded interface was “clean, not overwhelming” (S2-P1), “task-centered” (S2-P4), “lightweight and approachable” (S2-P7), and with “limited distractions” (S2-P8). S2-P2 also noted that, “I can imagine a less fighting the UI if I were a newbie,” indicating reduced cognitive friction during software navigation for beginners. Nevertheless, participants commented that the scaffolded interface may conflict with their familiarity with the default software layout. S2-P3 remarked that they “missed the old familiarity,” pointing to the tension when introducing novel interfaces.
Scaffolding’s impact on skill development (RQ4). Participants found that the scaffolded interface provided a supportive foundation for building practical skills, particularly for newcomers. For example, S2-P5 called it a “great stepping stone from beginner to intermediate,” and S2-P6 noted the interface made it “easier to build confidence in doing simple stuff.” These responses show that experts see our interface as helping beginner users build confidence and gain skills while learning to use professional creative software.
Using scaffolded interfaces in broader contexts (RQ5). Participants with experience in other professional creative software recognized the underlying principles of scaffolded interface design as being transferable. For instance, S2-P8 stated that, “I think the knowledge about motion, timing, bone constraints translates to using other tools.” Building on such transferability, participants suggested potential applications and impacts of using scaffolded interfaces in other software. S2-P6 said that, “3ds Max can benefit from this. The modifier stack alone is complex enough. If tools were grouped by task and revealed based on what you’re actually doing, it’d cut down on confusion.” Participants also anticipated the benefits, e.g., “it could help new users get productive faster, and even experienced users might find it useful to stay focused” (S2-P6). These results suggest user enthusiasm for applying the scaffolded interface generation method more broadly to reduce interface complexity and enhance workflow efficiency.
6.2.3. User Behavior
We observed that expert users approached tasks flexibly, sometimes bypassing the step-by-step workflow stages in the scaffolded interface. For instance, S2-P6 switched between marking seams (stage 1) and unwrapping & editing (stage 2). They started with the “Select Edge Loops” tool in stage 1 to mark seams, used “Unwrap” in stage 2 to check the UV map, then went back to use “Edge Select” in stage 1 to refine the seams and repeated the process. This reflects their preference for multitasking and iterative workflows. Furthermore, when needed tools were not presented at their current complexity level, several participants (S2-P2, S2-P5, S2-P6, S2-P8) sometimes did not look for these tools by advancing complexity levels. Instead, they accessed tools via keyboard shortcuts or the default Blender interface—the operations they were familiar with. This aligns with their feedback that the scaffolded interface was less supportive for tool exploration compared to the experience of beginners.
7. Discussion
Based on the user study findings, here we discuss the opportunities in AugGen, its limitations, and future directions.
7.1. Opportunities in AugGen
Task-aware adaptation to support software learners and educators. In our study, beginner users reported benefits in both task performance and concept learning with our scaffolded interface, showing its promise for instructional use. To better support learning for software learners, the interface can adapt its scaffolding strategy based on task complexity. For simple tasks, the interface can serve as a guided playground where future design could expand the available tools slightly beyond the current step to promote curiosity-driven exploration. For complex tasks, the focus shifts to practice and learning, so the interface can offer a smaller, essential toolset with clear links between domain concepts and software operations. It could also include best-practice tips or visual aids like video clips to help ensure key concepts and operations are grasped before moving on.
In addition, empowering educators is essential to unlock the full potential of using scaffolded interfaces for instruction. Educators could choose the concepts and tools the interface highlights, tailoring it to their teaching goals. Equally important is the control over the instruction process. If educators set when new tools and concepts appear, they can pace teaching based on student progress, creating a structured path that builds complexity gradually. By offering the blend of task-aware adaptation for learners and educators, scaffolded interfaces can evolve into helpful assistance for educational purposes.
Scaffolded interface driven by productivity and creativity. Our study with experts revealed that they focused on boosting productivity through personalized workflows and creativity by flexible exploration. Therefore, scaffolded interfaces targeting at expert end users should evolve beyond structured guidance towards systems driven by these needs. To support productivity, a scaffolded interface could learn a user’s unique shortcuts, tool sequences, and niche commands, and optimize workflows by rearranging palettes for top tools or suggesting more efficient alternatives. Implementing this requires robust user activity logging, potentially through screen monitoring by in-application AI copilots (cunningham2025copilotvision; google2025aistudio), to adapt interfaces according to the user’s demonstrated behavior.
To allow creative needs that step outside routine workflows, a scaffolded interface can act as a serendipity engine that encourages exploration. For instance, it can suggest unusual tool pairings based on deeper function relationships or highlight rarely used but relevant tools. This can be powered by LLMs, which use the user’s current state and project context to generate uncommon operations or workflows. We explore this potential by implementing scaffolded interface variations adapted to multiple creative demands, as shown in Appendix E.
Scaffolded interface meta-layer for cross-software workflows. An enthusiasm by our expert users was using the scaffolded interface generation method in multiple software platforms. Experts highlighted the transferable task-aware, concept-driven UI design for task efficiency, using the scaffolded interface as a meta-layer to bridge different applications. Imagine that end users could design their own scaffolded interface built upon the tools they frequently use. Instead of learning a new interface for every software, they could keep this consistent interface that maps familiar tools to different underlying software. Such a interface could help improve cross-software workflows as many complex projects require switching between multiple applications, and create a unified experience, reducing the effort required by switching contexts. In addition, this scaffolded interface meta-layer can help users stay focused on learning fundamental concepts and operations instead of spending significant efforts on learning each software’s new interface. The implementation of this idea would require technical efforts to map concepts and operations to different underlying software, but the potential to simplify cross-platform workflows can make it interesting to explore for future research.
7.2. Limitations and Future Directions
Our work has some limitations that we discuss below. First, our interface uses user-selectable complexity levels for progressive tool disclosure. This is just one point on the spectrum from user-driven to system-driven control. Future research could explore system-driven interface adaptation based on performance metrics or user modeling, as well as mixed-initiative approaches where the system suggests level changes or adapts specific interface elements contextually. Second, some expert participants raised concerns about potential over-reliance on the scaffolded interface, which might limit learning of the full software. To address this, future design could include cues like UI minimaps or short animations showing where tools come from. The scaffolded interface could also fade as users gain skill, encouraging further transition to the native UI. Third, expert feedback suggested that they desired UI customization for personalized workflows. This encourages the exploration of combining on-demand and scaffolded UI generation. Future research could use our technical pipeline, which allows interface iteration with human control, to let users customize their interfaces and evaluate such dynamic user experience. Lastly, our technical implementation demonstrated an example of building scaffolded interfaces in a software tool for 3D modeling and animation tasks. As introduced in Section 3, our approach can generate UIs for broader tasks and software that offers manuals and APIs. Future research could test the implementation of scaffolded interfaces in other applications and domains to explore broader use and impact.
8. Conclusion
Steep learning curves in professional creative software due to complex interfaces significantly hinder user learning and skill development. In this paper, we introduced AugGen, a method for generating scaffolded interfaces in professional creative software to support task-based learning. The scaffolded interfaces present task-relevant tools, manage complexity through adjustable levels, organize the UI around workflow stages with domain concepts integrated, and connect users to native software interactions to support skill transfer. Our evaluation involving both beginners and experts showed that the scaffolded interfaces significantly improved task performance for both user groups and supported learning during task execution for beginners. These findings demonstrated the effectiveness of combining task-awareness, workflow-based organization, concept integration, and progressive feature disclosure in interface design. Based on a discussion of the study results, we provided insights for future research on user empowerment through scaffolded interfaces for software learners, educators, and end users.
Appendix A LLM Prompt
A.1. Workflow Analysis and Concept Identification
You are an expert task workflow analyst specializing in task decomposition for a user task. Your objective is to analyze the user task and break it down into its primary workflow stages. TASK TO ANALYZE: {USER TASK DESCRIPTION} METHOD: Decompose the task into logical stages that represent a typical, efficient workflow a user would follow within {SOFTWARE}. Focus on the major distinct phases, moving from initial setup/preparation towards finalization/review. Aim for a reasonable number of stages that capture the core workflow without excessive granularity. Also, extract domain concepts relevant to each workflow stage. For each domain concepts, provide a concise, straightforward explanation for a human to easily understand. OUTPUT REQUIREMENTS: Provide the output as a numbered list. Each list item must represent a distinct workflow stage and include: A concise Stage Name. A brief, one-sentence Stage Description clarifying the purpose of the stage. Domain Concepts related to the stage. A concise, straightforward Concept Explanation explaining the concept. The output must be structured following the EXAMPLE OUTPUT FORMAT. EXAMPLE OUTPUT FORMAT: Stage Name 1: Stage Description, Domain Concepts, Concept Explanation. Stage Name 2: Stage Description, Domain Concepts, Concept Explanation. Stage Name n: Stage Description, Domain Concepts, Concept Explanation.
A.2. Tool Selection and Complexity Assessment
You are an expert user of {SOFTWARE}. Your task is to identify relevant tools in {SOFTWARE} to perform each workflow stage of {WORKFLOW STAGES}. WORKFLOW STAGES: {STAGE NAME}: {STAGE DESCRIPTION}, {DOMAIN CONCEPTS}, {CONCEPT EXPLANATION} METHOD: Identify the keyboard operators, mouse operations, and UI interactions needed to perform each workflow stage. The identified tools must be common and essential for the workflow stage. Prioritize stable and frequently used tools. In addition, you need to assign a difficulty level for each identified tool by referring to {SOFTWARE MANUALS}. The difficulty level should be basic, intermediate, or advanced for a human user. OUTPUT REQUIREMENTS: For each stage, output a bulleted list of selected tools. Each item in the list must include: The Stage Name from the {WORKFLOW STAGES}. The Selected Tools required to perform the task in the stage. A Rationale that briefly and clearly explain why the selected tools are essential and necessary. A Complexity Level of Basic, Intermediate, or Advanced for the selected tools. EXAMPLE OUTPUT FORMAT: Stage Name 1: Selected Tools, Rationale, Complexity Level. Stage Name 2: Selected Tools, Rationale, Complexity Level. Stage Name n: Selected Tools, Rationale, Complexity Level.
A.3. UI Code Generation and Tool Labeling
You are a coding expert for {SOFTWARE}. Your task is to generate the code to implement {SELECTED TOOLS} based on {SOFTWARE} functionality and generate UI code to display {SELECTED TOOLS} in a user interface for user interaction. SELECTED TOOLS: {STAGE NAME}: {SELECTED TOOLS}, {RATIONALE}, {COMPLEXITY LEVEL}. METHOD: For each selected tool in {SELECTED TOOLS}, you need to translate the tool operation into the corresponding code execution by referring to {SOFTWARE APIS}. You must make sure the code can achieve the identical result using the tool. The code you generate should include all the required parameters and properties for the code to be correctly executed. Next, you need to convert the code into interactive UI controls that can be used by a human. You can use buttons, dropdown menus, radio buttons, tabs, text field, switch toggles, or other UI controls that are supported by {SOFTWARE} and suitable for each tool. You must refer to {SOFTWARE APIS} to make sure the UI control can be implemented. To organize the code for UI controls, you should refer to the {EXAMPLE CODE} to group the code for each UI control under their {WORKFLOW STAGE} and {COMPLEXITY LEVEL}. In addition, you should implement a mouse-over tooltip for each UI control with their relevant {DOMAIN CONCEPT} and {CONCEPT DESCRIPTION}. EXAMPLE CODE WITH UI LAYOUT STRUCTURE {EXAMPLE CODE} OUTPUT REQUIREMENTS: Your output is a code script. The script should be standalone with all necessary imports, function definitions, and main function for it to be executable. You should refer to {EXAMPLE CODE} for the code script generation.
Appendix B User Study Materials
B.1. Walk Cycle Scaffolded UI
Figure 11 shows the scaffolded UI for the walk cycle task.
Figure 11. Walk cycle animation using the scaffolded UI.:
B.2. Study Guide
The video tutorial for UV unwrapping is available at https://www.youtube.com/watch?v=7JUNlj6mR0U, and for building a walk cycle can be found at https://www.youtube.com/watch?v=e_COc0ZVHr0.
B.3. Questionnaires
Table 1 summarizes the questions from NASA-TLX to evaluate task load. The custom questionnaires for task performance, concept learning, and user interface experience are presented in Tables 2, 3, and 4.
Table 1. Task load questionnaire (NASA-TLX).
Mental demand How mentally demanding was it using the interface to perform your task?
Temporal demand How hurried or rushed was it using the interface to perform your task?
Performance How successful were you in accomplishing what you were asked to do?
Effort How hard did you have to work to accomplish your level of performance?
Frustration How insecure, discouraged, irritated, stressed, and annoyed were you?
Table 2. Task performance questionnaire.
Workflow clarity I felt clear about the task workflow.
Task efficiency I was able to complete the task efficiently.
Tool relevance It was easy for me to find the tools I needed to perform my task.
User focus & engagement It was easy for me to stay focused on and engaged in my task.
User confidence & satisfaction I felt confident performing the task and was satisfied with my accomplishment.
Table 3. Concept learning questionnaire.
Concept-workflow correlation I was able to correlate task-related concepts with the task workflow.
Concept understanding Using this interface helped me to understand these task-related concepts.
Concept-tool correlation I was able to correlate tools used for my task with these task-related concepts.
Learning motivation Using this interface enhanced my motivation and experience of learning concepts.
Adjacent concept exploration I was able to explore concepts beyond those directly related to my task.
Table 4. User interface experience questionnaire.
Tool support vs. distraction The interface offered relevant tools that supported my task without unnecessary distractions.
Prior knowledge requirement Prior interface familiarity (e.g., tool locations, usage, keyboard shortcuts) was not required to perform my task.
External help requirement Extra assistance (e.g., referring to tutorials or asking questions) was not required to perform my task.
Adjacent tool exploration I was able to explore tools in the interface beyond those directly related to my task.
Interface vs. user adaptation The interface appeared tailored to support my task rather than requiring me to adapt to it.
B.4. Open-Ended Questions
Compared to the default Blender interface, please provide your experience with the custom interface. We would appreciate detailed responses.
1. Did the custom interface help you complete the task better or worse than you expected? Were you able to easily find what you needed in the interface to perform your task? How did it make you feel while working, e.g., did you feel engaged or sometimes lost? Did the design of the interface make sense to you?
2. (Task 1: UV Unwrapping) If you were to learn new concepts, would the custom interface help you understand the concepts better or make them more confusing? For example, would things like seams, unwrapping, and UV stretching be easier to grasp? Would it help you discover anything new that isn’t directly part of your task but a related idea such as UV islands? Would the way the interface presents the individual tools make learning feel natural or difficult?
2. (Task 2: Building Walk Cycle) If you were to learn new concepts, would the custom interface help you understand the concepts better or make them more confusing? For example, would things like key poses and bone constraints be easier to grasp? Would it help you discover anything new that isn’t directly part of your task but a related idea such as secondary motion? Would the way the interface presents the individual tools make learning feel natural or difficult?
3. How did the custom interface affect your overall experience? Did it feel like it was supporting you or getting in your way? Could you figure things out on your own, or did you need help? Did you feel like you had to adjust to the system, or did the system adjust to you and your needs?
4. If you were to use the custom interface for a longer period, would it make your long-term skill development easier or harder? For example, would building up your skills from basic to advanced be easier, and if so, how? Would it help you build practical skills by teaching you underlying concepts? Would the learned concepts be transferable?
5. If you have experience with other software similar to Blender (e.g., Maya, 3ds Max), what do you think about the impact of using similar custom interfaces in them? How might the custom interfaces be used?
6. Are there any features you would add, remove, or modify in the custom interface? If yes, please specify.
Appendix C User Study 2 Participant Demographic Information
Table 5 presents the demographic information of Study 2 participants.
Table 5. Study 2 participant demographic information.
ID Year of Experience 3D Modeling Animation Additional Software Experience
S2-P1 2 SketchUp
S2-P2 5 Maya, ZBrush, MotionBuilder
S2-P3 3 Maya
S2-P4 3 3ds Max, Mixamo, Unity, Unreal Engine
S2-P5 6 Unity, Maya
S2-P6 4 Tinkercad, 3ds Max, Maya
S2-P7 1 FreeCAD, Tinkercad
S2-P8 5 Substance 3D, After Effects
Appendix D User Study 1 ANOVA Results
Tables 6 to 9 report the two-way ANOVA results of Study 1.
Table 6. Two-way ANOVA results for task load measures.
Interface Effect Task Effect Interaction Effect
$F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$
Mental demand $22.53$ <.001<.001 .43.43 $8.53$ .007.007 .22.22 $0.13$ .718.718 .00.00
Temporal demand $95.38$ <.001<.001 .76.76 $4.71$ .039.039 .14.14 $0.52$ .475.475 .02.02
Performance $150.36$ <.001<.001 .83.83 $0.19$ .665.665 .01.01 $1.73$ .200.200 .05.05
Effort $69.96$ <.001<.001 .70.70 $8.51$ .007.007 .22.22 $8.51$ .007.007 .22.22
Frustration $91.43$ <.001<.001 .75.75 $4.32$ .047.047 .13.13 $2.77$ .107.107 .08.08
Table 7. Two-way ANOVA results for task performance measures.
Interface Effect Task Effect Interaction Effect
$F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$
Workflow clarity $34.60$ <.001<.001 .54.54 $2.05$ .164.164 .06.06 $0.63$ .433.433 .02.02
Task efficiency $19.96$ <.001<.001 .40.40 $12.55$ .001.001 .30.30 $0.21$ .648.648 .01.01
Tool relevance $48.13$ <.001<.001 .62.62 $0.53$ .471.471 .02.02 $0.13$ .718.718 .00.00
User focus & engagement $71.46$ <.001<.001 .70.70 $0.42$ .521.521 .01.01 $2.30$ .140.140 .07.07
User confidence & satisfaction $38.53$ <.001<.001 .56.56 $2.13$ .155.155 .07.07 $4.80$ .037.037 .14.14
Table 8. Two-way ANOVA results for concept learning measures.
Interface Effect Task Effect Interaction Effect
$F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$
Concept-workflow correlation $78.91$ <.001<.001 .73.73 $7.39$ .011.011 .20.20 $0.03$ .857.857 .00.00
Concept understanding $29.67$ <.001<.001 .50.50 $2.93$ .098.098 .09.09 $0.02$ .877.877 .00.00
Concept-tool correlation $33.39$ <.001<.001 .53.53 $1.98$ .171.171 .06.06 $1.20$ .284.284 .04.04
Learning motivation $36.25$ <.001<.001 .55.55 $0.49$ .490.490 .02.02 $0.18$ .678.678 .01.01
Adjacent concept exploration $41.20$ <.001<.001 .58.58 $10.30$ .003.003 .26.26 $0.16$ .691.691 .01.01
Table 9. Two-way ANOVA results for user interface experience measures.
Interface Effect Task Effect Interaction Effect
$F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$ $F_{(1,30)}$ $p$ $\eta^{2}_{p}$
Tool support vs. distraction $44.44$ <.001<.001 .60.60 $0.44$ .510.510 .01.01 $0.00$ $1.000$ .00.00
Prior knowledge requirement $32.36$ <.001<.001 .52.52 $11.65$ .002.002 .28.28 $2.14$ .155.155 .07.07
External help requirement $26.75$ <.001<.001 .47.47 $15.35$ .001.001 .34.34 $4.15$ .051.051 .12.12
Adjacent tool exploration $59.25$ <.001<.001 .66.66 $3.25$ .082.082 .10.10 $0.67$ .420.420 .02.02
Interface vs. user adaptation $104.14$ <.001<.001 .78.78 $0.14$ .708.708 .00.00 $1.29$ .266.266 .04.04
Appendix E Scaffolded Interface Variations
E.1. UV Unwrapping
Figure 12(a) shows the scaffolded interface variations for UV unwrapping. For organic models like characters and creatures, the workflow focuses on creating smooth, low-distortion texture maps. The interface helps users define seams to flatten curved surfaces, apply algorithms that preserve surface flow, and refine UV layouts to keep textures accurate during animation. For hard-surface models, the workflow is designed to handle their unique geometry and texturing needs. The interface supports multiple UV sets for different textures, uses precise projection and alignment tools to create clean UV islands, and includes tools for accurate decal placement.
E.2. Building Walk Cycle
Figure 12(b) presents the scaffolded interface variations for building a walk cycle. The emotive walk cycle workflow adds emotion and personality to walk animations. The interface guides users in analyzing reference material, adjusting key poses, especially spine and head, and refining timing and curves, with layered secondary motion to express the desired mood or personality. The stylized, abstract walk cycle workflow focuses on creating animations that move away from realism for artistic effect. The interface supports exaggerating form and motion, adjusting timing and interpolation for unique rhythms, and creating bold, non-realistic effects to match artistic or conceptual goals.
(a) UV unwrapping.
(b) Building walk cycle.
Figure 12. AugGen variations.: